find general and singular solution y=2xp+4yp^2;X=x,Y=y^2 ,by using clairat's equation


1
Expert's answer
2021-06-29T18:20:03-0400

"y=2xp+4yp^2"

Multiplying by y, we obtain

"y^2=x\cdot 2yp+4y^2p^2"

By using variables "X=x" , "Y=y^2" , "P=dY\/dX=2ydy\/dx=2yp" , we can re-write this equation as

"Y=XP+P^2"

This is Clairaut's equation. Differentiate it by X:

"dY\/dX=P=P+XP'+2PP'"

"P'(X+2P)=0"

The general solution: "P'=0" , "P=C",

"Y=XP+P^2=CX+C^2" ,

"y^2=Cx+C^2"

The singular solution: "X+2P=0" , "P=-X\/2" , "Y=XP+P^2=X(-X\/2)+(-X\/2)^2=-X^2\/4" ,

"y^2=-x^2\/4"

"y=\pm ix/2" (acceptable, if x, y take values in complex numbers).


Answer. The general solution: "y^2=Cx+C^2". The singular solution:

"y=\pm ix/2" (acceptable, if x, y take values in complex numbers).